5 Question 6 [10 points] Solve for z, and give your answer in the form a+bi. 2z-3-2i=z-5+5i z=0 Question 7 [10 points]

please respond asap
Answer #1

22-3-21 2 2-5751 z tw z žto »z=(3+2i) = 2-5+51 2z - 3721 = z-5759 let z=arbi zlathi) - (3-1) = atbilstsi zatzbe - (3-28) zatb